OpenStudy (xhawaiiangirlx):

Solve for x: 3x − 24 = 81

OpenStudy (anonymous):

x=35

OpenStudy (frozen15):

just asking, would the questions already be posted from other ppl or would I have to post?

OpenStudy (xhawaiiangirlx):

thanks both of you

OpenStudy (benlindquist):

3x = 105

OpenStudy (frozen15):

oops srry and Simplifying 3x + -24 = 81 Reorder the terms: -24 + 3x = 81 Solving -24 + 3x = 81 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '24' to each side of the equation. -24 + 24 + 3x = 81 + 24 Combine like terms: -24 + 24 = 0 0 + 3x = 81 + 24 3x = 81 + 24 Combine like terms: 81 + 24 = 105 3x = 105 Divide each side by '3'. x = 35 Simplifying x = 35

OpenStudy (benlindquist):

johnny, don't give out answers

OpenStudy (xhawaiiangirlx):

it's not 35

OpenStudy (frozen15):

I wasn't done

OpenStudy (xhawaiiangirlx):

thanks benlindquist

OpenStudy (xhawaiiangirlx):

and everyone else for effort

OpenStudy (benlindquist):

wy

OpenStudy (frozen15):

umm actually it is 35 , I have done this before and it was 35 so,,,

OpenStudy (xhawaiiangirlx):

It's not in the answer choice that's why

OpenStudy (xhawaiiangirlx):

When solving negative one over three.(x − 15) = −4, what is the correct sequence of operations? Multiply each side by -1/3 negative one over three , subtract 15 from each side Multiply each side by −3, add 15 to each side Multiply each side by - 1/3 negative one over three , add 15 to each side Multiply each side by 3, subtract 15 from each side

OpenStudy (frozen15):

oh ok my bad then .. well ur answer is 35 but 105 for the options